A steel company has three blast furnaces of varying sizes. If furnaces A,B, and C are used full time, 800 metric tons of steel produced per day. If A and B are used half time while C at full time, 545 metric tons are produced. If A is not used, B as full time, and C as half-time, 410 metric tons are produced. How many metric tons per day does each furnace produce?
Answer by ikleyn(52915)

.
Let a = "How many metric tons per day furnace A produces",
    b = same for B,
    c = same for C.

Then you have these equations:

a   + b   + c   = 800,     (1)
a/2 + b/2 + c   = 545,     (2)
      b   + c/2 = 410.     (3)

Now multiply eq(2) by 2 (both sides), and then subtract eq(1) from it. You will get

a + b +  c  = 800          (1)
a + b + 2c = 1090          (4)
--------------------subtract      eq(1) from eq(4)

         c = 1090 - 800 = 290.

Having this, find b from (3).

Having b and c, find a from (1).

For similar problems, see the lesson
    - Joint-work problems for 3 participants
in this site.


Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this textbook under the topic "Rate of work and joint work problems" of the section "Word problems".